# JMLR: Joint Medical LLM and Retrieval Training for Enhancing Reasoning and Professional Question Answering Capability

### v3

@@ 5.2 Joint Training > RAG
Llama-7B Llama-7B-ColBERT FT-Llama-7B FT-Llama-7B-ColBERT JMLR-7B-Seperate JMLR-7B
MedQA 24.98 24.01 44.00 40.60 55.30 56.20
Amboss 17.10 35.90 46.50 45.60 71.00 71.20
@@
During fine-tuning, synchronously updating both the Retriever and LLM yields better results than updating only the LLM. Table 2 shows that JLMR models significantly outperform traditional RAG methods. For example, JLMR-13B’s scores (72.8% on MMLU, 65.5% on MedMcQA) is far superior to RAG-13B (69.9% on MMLU, 64.2% on MedMcQA). This result suggests that without updating the retriever, the documents retrieved may be relevant to the question but not necessarily helpful to the LLM’s response. Joint training ensures that the retriever learns which documents are beneficial for Llama-2.
To evaluate the adaptability of the JLMR retriever, we further applied it to other LLMs. Specifically, we employed different retrievers to identify relevant documents, which were then provided to ChatGPT for generating the final answer. The retriever trained with JMLR attained an average accuracy of 57.5%, while the retriever from RAG achieved an accuracy of 56.7%. These results indicate that the JLMR retriever effectively identifies documents that not only support Llama but also aid ChatGPT in selecting the correct answers. Interestingly, this experimental outcome aligns with recent studies on weak-to-strong generalization Burns et al. (2023), such as the Aligner Ji et al. (2024). It appears that training a retrieval module using a weaker model through the JMLR process enables it to function as a model-agnostic plug-and-play module, which subsequently enhances the performance of a stronger model. This suggests a promising research direction for future studies.
6 Ablation Study
To assess the effects of fine-tuning and information retrieval on model performance, we structured our experiments as follows:
  • Original Model Evaluation: Test the original Llama 2-7B on the MedQA and Amboss datasets.
  • Retrieval Augmentation: Evaluate the addition of a retriever during the inference stage to assist Llama 2-7B (Llama-7B+Colbert) in answering questions, analyzing the impact of the Colbert retriever.
  • Fine-Tuning: Fine-tune Llama 2-7B on the Amboss and MedQA training sets to develop FT-Llama-7B.
  • Retrieval with Fine-Tuning: Add a retriever during the inference stage to assist the fine-tuned model (FT-Llama-7B+ColBERT), to explore the effects of Colbert on the fine-tuned model.
  • Asynchronous Training: Fix the parameters of the retriever while only updating those of Llama, then reverse to update only the retriever’s parameters using rank loss, resulting in JMLR-7B-Separate.
  • Joint Training: Train the retriever and Llama simultaneously (JMLR-7B) to compare against separate training and to examine the impact of joint training on enhancing model performance.
Table 4 displays the varied performance metrics of the models at different experimental stages. Observations are as follows:
  • Llama-7B+Retriever (24.01% on MedQA, 35.90% on Amboss) underperformed compared to the original Llama 2-7B (24.98% on MedQA, 17.10% on Amboss), suggesting that the original Llama 2-7B model struggles to effectively utilize external documents, with longer documents potentially leading to misunderstandings or erroneous information synthesis.
@@ 6 Ablation Study
  • The FT-Llama-7B (44.0% on MedQA, 46.5% on Amboss) showed significant improvements over Llama-7B, indicating that the fine-tuning phase notably enhanced Llama’s capabilities in medical QA. However, this phase did not effectively leverage medical guidelines.
@@
  • The JMLR-7B-Separate model achieved superior performance (55.3% on MedQA and 71.0% on Amboss), surpassing the FT-Llama-7B-ColBERT (40.6% on MedQA, 46.5% on Amboss). The distinctive fine-tuning approach of JMLR-7B-Separate, which integrates medical documents during the tuning phase, enables the model to utilize medical resources more effectively.
@@
  • Finally, JMLR-7B (56.2% on MedQA, 71.2% on Amboss) uses a joint training approach that facilitates the retrieval of highly relevant medical documents, enhancing the model’s ability to accurately answer questions by presenting similar case-based treatment plans for consideration.
As illustrated in Table 5, JMLR retrieves closely related cases, significantly improving the model’s decision-making process by offering comparable treatment options.
7 Discussion: JMLR Effectiveness of the reasoning ability
To verify that our model can accurately answer questions and provide detailed reasoning, we tasked JMLR-13B with generating reasons for its responses to the Amboss and MedQA test sets. Similarly, we had ChatGPT generate corresponding reasons. We employed two methods for scoring these reasons: 1) We used GPT-4 to judge which model’s generated reasoning was better based on the ground truth. 2) We engaged three medical professionals to review and assess a sample of 20 questions drawn from the test sets. These doctors were asked to judge which model’s reasoning was closer to the actual, clinically accurate reasons.
In the assessment using GPT-4, JMLR-13B achieved a win rate of 0.63, while under the evaluation of medical professionals, our model attained a win rate of 0.60. These results suggest that simply increasing the model’s parameters is not the most effective way to expand medical knowledge. Instead, JMLR-13B’s ability to consult different medical guidelines for various questions and to find corresponding answers and explanations significantly reduces hallucinations and greatly enhances the model’s explanatory power. This achievement highlights the model’s proficiency in accessing and applying relevant medical information contextually, thereby providing more accurate and credible explanations. The close alignment of the win rates from both GPT-4 and medical professionals’ evaluations further underscores the model’s robustness and reliability in medical question-answering and reasoning tasks.
@@ 8 Related Work
In the clinical domain, LLMs’ factual inaccuracies due to gaps in medical knowledge pose significant risks, including misdiagnoses Petroni et al. (2019); Sung et al. (2021); Yao et al. (2022); Singhal et al. (2023a). Incorporating diverse knowledge repositories into LLMs enhances their performance across various NLP tasks, particularly in language modeling (Min et al., 2022). In traditional RAG models, the retrieval component is static, with updates applied only to the LLM. The process involves using the input as a query to first retrieve a collection of documents from a designated corpus, which the LLM then uses to enhance its predictions (Yu, 2022; Izacard et al., 2022). However, integrating retrieved documents does not guarantee effective use of external knowledge. Innovations like Atlas (Izacard et al., 2022) and RETRO (Borgeaud et al., 2022b) have modified models to better utilize retrieved information, but challenges remain in coherently blending this external data with the LLM’s pre-trained responses. Although some researchers have explored unsupervised training of retrievers to improve document relevancy for LLMs, such as REALM (Guu et al., 2020), REPLUG (Shi et al., 2023), and RA-DIT (Lin et al., 2023), these systems are usually trained independently from LLMs. Our approach innovatively integrates and synchronously trains both the retriever and the LLM, ensuring the retriever supplies more appropriate documents, thereby boosting the LLM’s effectiveness. We have validated our method’s superiority through comparisons with step-by-step training techniques, confirming its enhanced performance in seamlessly integrating external knowledge.
@@ A.1 Case Study
We presented four specific examples to further understand why JLMR outperforms other public models. The first three examples are from the Amboss dataset. Due to privacy concerns with Amboss, we cannot display the complete content of these questions. The last example is from a USMLE question. Our answers for the first two questions are correct, while ChatGPT responded incorrectly. Although both our model and ChatGPT answered the last two questions correctly, there were some differences in the reasoning generated. For the first example, we found that JLMR’s retrieval function could extract similar questions, leading the model to answer incorrectly regarding direction. The retrieved document served as a background, fully utilizing the previously fine-tuned dataset. In the second example, we noticed that ChatGPT made a common-sense error: vaccination is not recommended for pregnant women in such cases. Even though ChatGPT has learned far more medical documents than our model by using pretraining or finetuning, it does not mean that the model will apply all medical knowledge correctly, even for some basic medical issues. In the third example, both we and ChatGPT answered correctly, but the reasoning generated by ChatGPT differed significantly from the correct answer’s reasoning: ChatGPT simply stated some facts without providing a logical reason. However, our model gave a more detailed and logical explanation of why there are decreased circulating T cells, as JLMR could retrieve related documents, such as some documents about SCID, to better understand the underlying principles. The last example differs from the first three, as the highest-scoring document retrieved was from public guideline. The guideline provides a detailed introduction to PCOS and its symptoms, helping the model better explain. In contrast, ChatGPT simply correlated the symptoms straightforwardly without thoroughly analyzing other related symptoms, which can often lead to diagnostic errors, even though it answered this question correctly. Our model, however, performed a comprehensive analysis.
@@ A.2 Accuracy Variation with Different Numbers of Documents
During training, we set the number of background documents to seven, maintaining this count during the inference phase as well. This quantity is optimal, according to our experiments, as illustrated in the figure 3. We employed the JLMR method to train the LLM on the MedQA dataset, retrieving varying numbers of documents. The performance is at its weakest when only one document is retrieved; this insufficiency leads to a lack of adequate medical knowledge for the model. As we increase the number of retrieved documents, the model’s performance gradually improves, reaching its peak with seven documents. However, once this number increases to ten, performance again declines due to retrieving an excess of irrelevant documents, which hinders the model’s ability to answer questions effectively.
Figure 3: The horizontal axis represents the number of documents retrieved in JLMR, while the vertical axis shows the accuracy of JLMR tested on the USMLE.
@@ 1 Q1: A 48-year-old man…
ChatGPT: The correct answer is A) Cerebral edema. Explanation: This patient presents with symptoms consistent with alcohol withdrawal syndrome (AWS), which can manifest as seizures (alcohol withdrawal seizures) due to abrupt cessation or reduction of alcohol intake in a chronic heavy drinker. AWS seizures typically occur within 48 hours of reducing alcohol intake. This patient’s laboratory findings are indicative of hyponatremia (low serum sodium), which is common in severe alcohol withdrawal due to the syndrome of inappropriate antidiuretic hormone secretion (SIADH). SIADH leads to water retention and dilutional hyponatremia. The risk of cerebral edema is increased in patients with severe hyponatremia who are rapidly corrected to normal serum sodium levels. Rapid correction of hyponatremia can lead to osmotic shifts of water into the brain cells, causing cerebral edema, which may result in neurological deterioration, seizures, coma, and death. This risk is particularly high in chronic alcoholics due to adaptive changes in brain osmolytes. Given the patient’s current status, urgent treatment to normalize serum sodium levels (such as with hypertonic saline infusion) puts him at increased risk for cerebral edema. Therefore, careful monitoring and slow correction of hyponatremia are essential in this patient to avoid this potentially fatal complication.
@@
Expertise Comment: ChatGPT incorrectly interpreted the patient’s symptoms as primarily related to Alcohol Withdrawal Syndrome (AWS) without adequately considering the patient’s laboratory results, particularly the severe hyponatremia (serum sodium 112 mEq/L). Although seizures can indeed accompany AWS, in this case, the diagnosis and treatment of hyponatremia should be the primary focus. Moreover, it incorrectly associated the risks of rapid correction of hyponatremia with cerebral edema. In fact, the main risk of rapidly correcting hyponatremia is osmotic myelinolysis, not cerebral edema. Osmotic myelinolysis results from rapid changes in blood osmolarity, leading to damage to the myelin sheath of brain neurons.
@@
ChatGPT: Based on the clinical presentation described, the most likely diagnosis for this patient is Rubella (German measles) infection during pregnancy, which can lead to congenital rubella syndrome (CRS) in the fetus if contracted during the first trimester. The typical symptoms of rubella include a maculopapular rash starting on the face and spreading downwards, along with lymphadenopathy and joint swelling. Given the risk of congenital rubella syndrome to the fetus, the most appropriate next step in management is: C) Vaccination Vaccination against rubella is contraindicated during pregnancy due to the risk of causing CRS. However, after delivery, the patient should receive the rubella vaccine to prevent future infections. This will not only protect the patient from future rubella infections but also prevent the risk of transmitting the virus to others, including vulnerable populations such as pregnant women. Options A, D, and E are not appropriate in this scenario. Oral acyclovir therapy (Option A) is used for herpes simplex virus infections, not rubella. Reassurance and symptomatic therapy (Option D) may provide comfort but do not address the potential risk to the fetus. Intravenous immunoglobulin therapy (Option E) is not indicated for rubella infection during pregnancy. Termination of pregnancy (Option B) is not necessary unless there are severe fetal abnormalities detected on further evaluation.
@@
Expertise Comment: ChatGPT made a made a basic error; it’s not appropriate to vaccinate during pregnancy.
Table 5: Case Study For Retrieval Documents
@@ 1 Q4: A 1-year-old girl is brought to the physician…
ChatGPT: Given the clinical presentation and findings described, this patient is most likely suffering from severe combined immunodeficiency (SCID), which is characterized by a deficiency in both T and B lymphocytes. The additional finding most likely to be present in this patient with SCID is: D) Decreased circulating T cells SCID is a primary immunodeficiency disorder characterized by profound defects in both cellular and humoral immunity. Patients with SCID typically have absent or markedly reduced T cell counts, as T cell development and function are severely impaired. This deficiency in T cells predisposes patients to recurrent, severe, and often life-threatening infections. While B cells may also be affected in SCID, the absence of T cells is the primary hallmark of the disease. Therefore, decreased circulating T cells (Option D) would be the most likely additional finding in this patient. Other options such as increased circulating IgE (Option B) or increased circulating neutrophils (Option F) are less specific to SCID and could be seen in a variety of other conditions.
@@
Expertise Comment: ChatGPT’s explanation of the pathology is unclear, such as the specific type of SCID (Severe Combined Immunodeficiency caused by adenosine deaminase deficiency) and how this condition affects the proliferation and maturation of lymphocytes, leading to a reduced number of T cells.
@@
ChatGPT: The most likely diagnosis for this patient is: D: Polycystic ovarian syndrome (PCOS) PCOS is a common endocrine disorder in women of reproductive age, characterized by hyperandrogenism, irregular menstrual cycles, and polycystic ovaries on ultrasound. The clinical features seen in this patient, including acne, unwanted facial hair (hirsutism), obesity, irregular menstrual cycles, and a family history of type 2 diabetes mellitus, are all consistent with PCOS. Other options: A: Hypothyroidism - While hypothyroidism can present with weight gain and menstrual irregularities, it is less likely given the presence of hirsutism and obesity in this patient. B: Idiopathic hirsutism - This is unlikely as idiopathic hirsutism typically presents with hirsutism without other features of PCOS such as irregular menstrual cycles and obesity. C: Cushing’s disease - While Cushing’s disease can present with weight gain and hirsutism, it is less likely in this patient as there are no signs of hypercortisolism such as central obesity, buffalo hump, or purple striae. E: Ovarian hyperthecosis - This is a rare condition characterized by luteinization of ovarian stromal cells leading to hyperandrogenism. However, PCOS is a more common cause of hirsutism and irregular menstrual cycles in young women.
@@
Expertise Comment: Our explanation is slightly more detailed in terms of linking the patient’s specific symptoms to the characteristics of PCOS and explaining why other diagnoses are less likely. It provides a more comprehensive understanding of PCOS and its relation to the patient’s symptoms and test results. On the other hand, ChatGPT offers a more concise but still accurate diagnosis.

### v4

@@ 5.2 Joint Training > RAG
During fine-tuning, synchronously updating both the Retriever and LLM yields better results than updating only the LLM. Table 2 shows that JMLR models significantly outperform traditional RAG methods. For example, JMLR-13B’s scores (72.8% on MMLU, 65.5% on MedMcQA) is far superior to RAG-13B (69.9% on MMLU, 64.2% on MedMcQA). This result suggests that without updating the retriever, the documents retrieved may be relevant to the question but not necessarily helpful to the LLM’s response. Joint training ensures that the retriever learns which documents are beneficial for LLM. To evaluate the adaptability of the JMLR retriever, we further applied it to other LLMs. Specifically, we employed different retrievers to identify relevant documents, which were then provided to GPT-3.5 for generating the final answer. The retriever trained with JMLR attained an average accuracy of 57.5%, while the retriever from RAG achieved an accuracy of 56.7%. In addition, we present a case study in Table 6, where JMLR retrieves closely related cases. These results indicate that the JMLR retriever effectively identifies documents that not only support Llama but also aid GPT-3.5 in selecting the correct answers.
5.3 Enhancing Model Rationale Quality
Haiku Opus GPT3.5 JMLR
UMLS Factuality 0.2337 0.2356 0.2187 0.2463
GPT-4 Overall 4.0559 4.2449 4.0620 4.3036
GPT-4 Comprehension 4.7107 4.7561 4.7549 4.7661
GPT-4 Reasoning 3.6701 4.0976 3.6863 4.0101
GPT-4 Recall 3.6726 3.8780 3.6849 4.1661
Table 4: UMLS-F and GPT-4 score across 4 different models (e.g., Claude3, GPT3.5, JMLR-13B).
In addition to significantly improving the model’s performance in QA accuracy, JMLR also plays a crucial role in helping the model generate higher-quality rationales with fewer hallucinations. As shown in the Table 4, both UMLS Factuality and GPT-4 Overall scores show that JMLR-13B outperforms the other three models, demonstrating its high-quality rational generation. This result aligns with our expert evaluations on the small samples, where JMLR-13B achieved a win rate of 0.60 compared to GPT-3.5.
Specifically, GPT-4 Comprehension primarily measures the plausibility of generated reasons. As shown in Table 4, the scores for GPT-4 Comprehension are uniformly high across all models, with only minor differences, indicating that all these models can generate logically coherent explanations for the questions. GPT-4 Reasoning evaluates medical logical reasoning abilities. JMLR-13B stands out in this critiria because its retrieval guidelines include valuable content, such as information on Differential Diagnosis, which aids the model in more effectively associating key information like symptoms, diseases, and medications presented in the questions. Finally, GPT-4 Recall assesses factuality, i.e., whether the model can hit the critical information points. Our results demonstrate that JMLR-13B achieves the most significant improvement in this criteria. The high-quality documents retrieved by JMLR-13B provide the necessary key information, allowing the model to generate higher-quality rationales around these key points. The assessments of UMLS Factuality are similar to GPT-4 Recall, and their final results are consistent. We further explore this in the appendix A and provide case studies to support our findings.
6 Ablation Study
MedQA Amboss
FT-Llama-7B 44.0 46.5
FT-Llama-7B-ColBERT 40.6 45.6
JMLR-7B-Separate 55.3 69.0
JMLR-7B 56.2 71.2
@@ 6 Ablation Study
To assess the effects of fine-tuning and information retrieval on QA accuracy, we compare JMLR-7B (on medical documents) with the following new baselines: The JMLR-7B-Separate model updates the parameters of both Llama and the retriever similar to JMLR-7B, but in an asynchronous manner. Initially, it fixes the parameters of the retriever while only updating those of Llama. Subsequently, it reverses this process, updating only the retriever’s parameters using rank loss. The FT-Llama-7B+ColBERT model, on the other hand, freezes the retriever in a manner akin to RAG-7B. However, the retriever in this setup is solely applied during inference and not during fine-tuning. The FT-Llama-7B serves as a naive fine-tuned baseline where the retriever is not employed. We fine-tuned the Llama 2-7B on the training sets. The accuracy are shown in Table 5. From the table, we can observe that:
@@
  • FT-Llama-7B+ColBERT (44.0% on MedQA, 46.5% on Amboss) showed significant improvements over FT-Llama-7B, indicating that the fine-tuning phase notably enhanced Llama’s capabilities in medical QA. However, this phase did not effectively leverage medical guidelines.
@@
  • JMLR-7B-Separate model achieved superior performance (55.3% on MedQA and 69.0% on Amboss), surpassing the FT-Llama-7B-ColBERT (40.6% on MedQA, 45.6% on Amboss). The distinctive fine-tuning approach of JMLR-7B-Separate, which integrates medical documents during the tuning phase, enables the model to utilize medical resources more effectively.
@@
  • Finally, JMLR-7B (56.2% on MedQA, 71.2% on Amboss) uses a joint training approach that facilitates the retrieval of highly relevant medical documents, surpassing the JMLR-7B-Seperate (55.3% on MedQA, 69.0% on Amboss). The result demonstrate that jointly training Llama and ColBERT outperforms training the two separately.
@@ 7 Related Work
In the clinical domain, LLMs’ factual inaccuracies due to gaps in medical knowledge pose significant risks, including misdiagnoses Petroni et al. (2019); Sung et al. (2021); Yao et al. (2022); Singhal et al. (2023a). Incorporating diverse knowledge repositories into LLMs enhances their performance across various NLP tasks, particularly in language modeling (Min et al., 2022). In traditional RAG models, the retrieval component is static, with updates applied only to the LLM. The process involves using the input as a query to first retrieve a collection of documents from a designated corpus, which the LLM then uses to enhance its predictions (Yu, 2022; Izacard et al., 2022). However, integrating retrieved documents does not guarantee effective use of external knowledge. Innovations like Atlas (Izacard et al., 2022) and RETRO (Borgeaud et al., 2022b) have modified models to better utilize retrieved information, but challenges remain in coherently blending this external data with the LLM’s pre-trained responses. Although some researchers have explored unsupervised training of retrievers to improve document relevancy for LLMs, such as REALM (Guu et al., 2020), REPLUG (Shi et al., 2023), and RA-DIT (Lin et al., 2023), these systems are usually trained independently from LLMs. Our approach innovatively integrates and synchronously trains both the retriever and the LLM, ensuring the retriever supplies more appropriate documents, thereby boosting the LLM’s effectiveness. We have validated our method’s superiority through comparisons with step-by-step training techniques, confirming its enhanced performance in integrating external knowledge.
@@ A.1 Case Study
We presented four specific examples to further understand why JMLR outperforms other public models as shown in Table 6 and Table  7. The first three examples are from the Amboss dataset. Due to privacy concerns with Amboss, we cannot display the complete content of these questions. The last example is from a USMLE question. Our answers for the first two questions are correct, while GPT-3.5 responded incorrectly. Although both our model and GPT-3.5 answered the last two questions correctly, there were some differences in the rationale generated. For the first example, we found that JMLR’s retrieval function could extract similar questions, leading the model to answer incorrectly regarding direction. The retrieved document served as a background, fully utilizing the previously fine-tuned dataset. In the second example, we noticed that GPT-3.5 made a common-sense error: vaccination is not recommended for pregnant women in such cases. Even though GPT-3.5 has learned far more medical documents than our model by using pretraining or finetuning, it does not mean that the model will apply all medical knowledge correctly, even for some basic medical issues. In the third example, both we and GPT-3.5 answered correctly, but the rationale generated by GPT-3.5 differed significantly from the correct answer’s rationale: GPT-3.5 simply stated some facts without providing a logical reason. However, our model gave a more detailed and logical explanation of why there are decreased circulating T cells, as JMLR could retrieve related documents, such as some documents about SCID, to better understand the underlying principles. The last example differs from the first three, as the highest-scoring document retrieved was from public guideline. The guideline provides a detailed introduction to PCOS and its symptoms, helping the model better explain. In contrast, GPT-3.5 simply correlated the symptoms straightforwardly without thoroughly analyzing other related symptoms, which can often lead to diagnostic errors, even though it answered this question correctly. Our model, however, performed a comprehensive analysis.
@@ A.2 Accuracy Variation with Different Numbers of Documents
During training, we set the number of background documents to seven, maintaining this count during the inference phase as well. This quantity is optimal, according to our experiments, as illustrated in the figure 3. We employed the JMLR method to train the LLM on the MedQA dataset, retrieving varying numbers of documents. The performance is at its weakest when only one document is retrieved; this insufficiency leads to a lack of adequate medical knowledge for the model. As we increase the number of retrieved documents, the model’s performance gradually improves, reaching its peak with seven documents. However, once this number increases to ten, performance again declines due to retrieving an excess of irrelevant documents, which hinders the model’s ability to answer questions effectively.
Figure 3: The horizontal axis represents the number of documents retrieved in JMLR, while the vertical axis shows the accuracy of JMLR tested on the USMLE.
@@ 1 Q1: A 48-year-old man…
GPT-3.5: The correct answer is A) Cerebral edema. Explanation: This patient presents with symptoms consistent with alcohol withdrawal syndrome (AWS), which can manifest as seizures (alcohol withdrawal seizures) due to abrupt cessation or reduction of alcohol intake in a chronic heavy drinker. AWS seizures typically occur within 48 hours of reducing alcohol intake. This patient’s laboratory findings are indicative of hyponatremia (low serum sodium), which is common in severe alcohol withdrawal due to the syndrome of inappropriate antidiuretic hormone secretion (SIADH). SIADH leads to water retention and dilutional hyponatremia. The risk of cerebral edema is increased in patients with severe hyponatremia who are rapidly corrected to normal serum sodium levels. Rapid correction of hyponatremia can lead to osmotic shifts of water into the brain cells, causing cerebral edema, which may result in neurological deterioration, seizures, coma, and death. This risk is particularly high in chronic alcoholics due to adaptive changes in brain osmolytes. Given the patient’s current status, urgent treatment to normalize serum sodium levels (such as with hypertonic saline infusion) puts him at increased risk for cerebral edema. Therefore, careful monitoring and slow correction of hyponatremia are essential in this patient to avoid this potentially fatal complication.
@@
Expertise Comment: GPT-3.5 incorrectly interpreted the patient’s symptoms as primarily related to Alcohol Withdrawal Syndrome (AWS) without adequately considering the patient’s laboratory results, particularly the severe hyponatremia (serum sodium 112 mEq/L). Although seizures can indeed accompany AWS, in this case, the diagnosis and treatment of hyponatremia should be the primary focus. Moreover, it incorrectly associated the risks of rapid correction of hyponatremia with cerebral edema. In fact, the main risk of rapidly correcting hyponatremia is osmotic myelinolysis, not cerebral edema. Osmotic myelinolysis results from rapid changes in blood osmolarity, leading to damage to the myelin sheath of brain neurons.
@@
GPT-3.5: Based on the clinical presentation described, the most likely diagnosis for this patient is Rubella (German measles) infection during pregnancy, which can lead to congenital rubella syndrome (CRS) in the fetus if contracted during the first trimester. The typical symptoms of rubella include a maculopapular rash starting on the face and spreading downwards, along with lymphadenopathy and joint swelling. Given the risk of congenital rubella syndrome to the fetus, the most appropriate next step in management is: C) Vaccination Vaccination against rubella is contraindicated during pregnancy due to the risk of causing CRS. However, after delivery, the patient should receive the rubella vaccine to prevent future infections. This will not only protect the patient from future rubella infections but also prevent the risk of transmitting the virus to others, including vulnerable populations such as pregnant women. Options A, D, and E are not appropriate in this scenario. Oral acyclovir therapy (Option A) is used for herpes simplex virus infections, not rubella. Reassurance and symptomatic therapy (Option D) may provide comfort but do not address the potential risk to the fetus. Intravenous immunoglobulin therapy (Option E) is not indicated for rubella infection during pregnancy. Termination of pregnancy (Option B) is not necessary unless there are severe fetal abnormalities detected on further evaluation.
@@
Expertise Comment: GPT-3.5 made a made a basic error; it’s not appropriate to vaccinate during pregnancy.
Table 6: Case Study For Retrieval Documents.
@@ 1 Q4: A 1-year-old girl is brought to the physician…
GPT-3.5: Given the clinical presentation and findings described, this patient is most likely suffering from severe combined immunodeficiency (SCID), which is characterized by a deficiency in both T and B lymphocytes. The additional finding most likely to be present in this patient with SCID is: D) Decreased circulating T cells SCID is a primary immunodeficiency disorder characterized by profound defects in both cellular and humoral immunity. Patients with SCID typically have absent or markedly reduced T cell counts, as T cell development and function are severely impaired. This deficiency in T cells predisposes patients to recurrent, severe, and often life-threatening infections. While B cells may also be affected in SCID, the absence of T cells is the primary hallmark of the disease. Therefore, decreased circulating T cells (Option D) would be the most likely additional finding in this patient. Other options such as increased circulating IgE (Option B) or increased circulating neutrophils (Option F) are less specific to SCID and could be seen in a variety of other conditions.
@@
Expertise Comment: GPT-3.5’s explanation of the pathology is unclear, such as the specific type of SCID (Severe Combined Immunodeficiency caused by adenosine deaminase deficiency) and how this condition affects the proliferation and maturation of lymphocytes, leading to a reduced number of T cells.
@@
GPT-3.5: The most likely diagnosis for this patient is: D: Polycystic ovarian syndrome (PCOS) PCOS is a common endocrine disorder in women of reproductive age, characterized by hyperandrogenism, irregular menstrual cycles, and polycystic ovaries on ultrasound. The clinical features seen in this patient, including acne, unwanted facial hair (hirsutism), obesity, irregular menstrual cycles, and a family history of type 2 diabetes mellitus, are all consistent with PCOS. Other options: A: Hypothyroidism - While hypothyroidism can present with weight gain and menstrual irregularities, it is less likely given the presence of hirsutism and obesity in this patient. B: Idiopathic hirsutism - This is unlikely as idiopathic hirsutism typically presents with hirsutism without other features of PCOS such as irregular menstrual cycles and obesity. C: Cushing’s disease - While Cushing’s disease can present with weight gain and hirsutism, it is less likely in this patient as there are no signs of hypercortisolism such as central obesity, buffalo hump, or purple striae. E: Ovarian hyperthecosis - This is a rare condition characterized by luteinization of ovarian stromal cells leading to hyperandrogenism. However, PCOS is a more common cause of hirsutism and irregular menstrual cycles in young women.
@@
Expertise Comment: Our explanation is slightly more detailed in terms of linking the patient’s specific symptoms to the characteristics of PCOS and explaining why other diagnoses are less likely. It provides a more comprehensive understanding of PCOS and its relation to the patient’s symptoms and test results. On the other hand, GPT-3.5 offers a more concise but still accurate diagnosis.
A.3 Training Resource
In our study, we utilized various training resources to evaluate the performance of different models on medical question-answering (QA) datasets. According to the data, for MedQA datasets, JMLR-13B, which requires 148 GPU hours for training, significantly outperforms Meditron-70B, which necessitates 42,630 GPU hours. In comparison, JMLR-7B needs approximately 100 GPU hours, RAG-7B requires 72 GPU hours, and RAG-13B needs 128 GPU hours.
The training time for JMLR models is considerably lower than that for Meditron-70B, though slightly higher than RAG-13B. However, JMLR models achieve state-of-the-art (SOTA) performance. Interestingly, we observe a substantial performance improvement in RAG-13B compared to RAG-7B. This improvement can be attributed to the fact that RAG models struggle to effectively learn how to understand or utilize medical documents. As a result, the performance of RAG-7B is limited by its model size. Conversely, RAG-13B, with its increased model size, is better able to comprehend documents, leading to enhanced performance. Our JMLR models, on the other hand, excel in enabling the model to understand and utilize medical documents effectively. This capability is reflected in their SOTA performance across multiple datasets, demonstrating the efficacy of our approach in medical QA tasks.
A.4 Factuality metrics: UMLS-F1
The assessment of factual accuracy in LLMs output leverages the UMLS concept overlap metric. The Unified Medical Language System (UMLS), established by Bodenreider (2004), significantly contributes to the biomedical domain’s interoperability. It achieves this by amalgamating and disseminating a comprehensive collection of biomedical terminologies, classification systems, and coding standards from many sources. By doing so, UMLS aids in reconciling semantic variances and representational disparities found across different biomedical concept repositories.
For the identification and alignment of medical named entities within texts to their corresponding biomedical concepts in UMLS, we employed the Scispacy library  [^3] We used the Scispacy en_core_sci_lg model.. Scispacy excels in identifying and clarifying entities, thus facilitating the accurate association of named entities found in LLMs output with the relevant UMLS concepts. This capability is critical for evaluating the LLMs output’s factual accuracy.
The analytical process for LLMs output utilizes metrics of precision and recall. Precision represents the ratio of concepts present in both the LLM output and ground truth content, serving as a measure of the LLM output’s factual correctness. In contrast, recall evaluates how well the information in the LLM output matches the intended content, reflecting the relevance of the presented information.
To calculate these metrics, we consider the concept sets from both the ground truth ($C_{ref}$) and the LLM output ($C_{gen}$). The formulas for recall and precision are as follows:
$\text{Recall}=\frac{|C_{ref}\cap C_{gen}|}{|C_{ref}|}$
$\text{Precision}=\frac{|C_{ref}\cap C_{gen}|}{|C_{gen}|}.$
The F1 score, derived from the above precision and recall values, is reported to provide a balanced measure of LLMs output’s accuracy and relevance.
A.5 Prompt for Evaluation
For GPT-4 evaluation, we assessment each result from three ways: comprehension, reasoning step and recall of knowledge. Here is our prompt:
GPT-4 System
Act as a USMLE evaluator, your role involves assessing and comparing a medical student’s explanation to the provided target answer. Begin the assessment by carefully reviewing the provided target answer. Then, based on following specific criteria, determine the score for the student’s answer.
Evaluation Criteria
For each diagnosis, evaluate the medical student explanation base on the following three questions:
Question 1
Does the medical student’s answer contain any evidence of incorrect reading comprehension? (indication the question has not been understood)
Question 2
Does the medical student’s answer contain any evidence of incorrect reasoning steps? (incorrect rationale for answering the question)
Question 3
Does the medical student’s answer contain any evidence of incorrect recall of knowledge? (mention of an irrelevant and/or incorrect fact for answering the question)
InputMedical student’s answer: Reason generated by model
Medical student’s answer: Reason generated by model
Target answer: Ground truth reason
Background Question: Question
Output Format Your evaluation should be provided in JSON format, as follows(don’t generate any other information): {{"case 1": {{"question 1": "The score for question 1", "question 2": "The score for question 2", "question 3": "The score for question 3", }}, "case 2": "score for case 2 with the same format as case 1","case 3": "score for case 3 with the same format as case 1", "overall score": "the average score for question 1, 2, 3", "reason": "the reason why you give the score"}}
Table 8: Evaluation Prompt for GPT-4.